The current in a series circuit of R = 50 and L= 50 mH lags the applied voltage by 80 °,the angular frequency w is:


 Select one:

 a. 945 rad/s

 b. 473 rad/s

 c. 709 rad/s

 d. 405 rad/s

 e. 567 rad/s


1
Expert's answer
2021-06-09T14:17:01-0400
"\omega=\frac{R}{L}\\\omega=\frac{50}{0.05}=1000\frac{rad}{s}"